I need the answer to this equation (6 ^ 10 ^ 4) / (6 ^ 10 ^ 3) Thanks
Answer by user_dude2008(1862)

(6 ^ 10 ^ 4) / (6 ^ 10 ^ 3) = 6^(10 ^ 4-10 ^ 3) = 6^(10^3(10-1)) = 6^(9*10^3)